The quotient is 712, the remainder is 36 and the divisor is 9 more than the sum of both. What is the dividend
Answer by ikleyn(52780)

.
The sum of both is  712 + 36 = 748.


The divisor is  748 + 9 = 757.


The dividend is  712*757+36 = 539020.